If Santa Claus ๐ŸŽ… distributes gifts to every single child in every nook and cranny of the world in a single night, (taking every unit surface area of the earth to contain a single child) and assuming he has the entire globe to distribute to, in only a matter of 8.5 hours, what is his speed? [ std. value of earth's radius is considered ]
โ€‹
โ€‹
โ€‹

Grade:12th pass

1 Answers

Manas Baul
8 years ago
Radius of Earth= 6378.1 Km
Total Surface Area = 4/3\pi r^3ย =ย 510,072,000 km2
Every Unit Surface Area Has one childย 
\thereforeย Total Number of Childs =ย 510,072,000
Total Time = 8.5 hrs
Hence, Speed =ย 510,072,000(2)(17)-1= 6.00084706 X 108ย Km/h
Which is way greater then speed of light.ย